OpenStudy (anonymous):

how do u salve x2-6x=27

OpenStudy (anonymous):

simple

OpenStudy (imstuck):

You need to get it into standard form which is \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

yup

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm x^2-6x=27 }\) \(\large\color{midnightblue}{ \rm x^2-6x+9=27 +9 }\) \(\large\color{midnightblue}{ \rm (x-3)^2=36 }\) \(\large\color{midnightblue}{ \rm x-3=± 6 }\) \(\large\color{midnightblue}{ \rm x=3± 6 }\)

OpenStudy (imstuck):

Actually you can just subtract the 27 from both sides instead of completing the square. It might be easier to explain to someone who doesn't know the ins and outs of factoring.